OpenStudy (anonymous):

Verify the identity. cos 4x + cos 2x = 2 – 2 sin^2 2x – 2 sin^2 x

OpenStudy (raden):

cos4x = cos2(2x) = 1 - 2sin^2 2x cos2x = 1 - 2sin^2 x just joining the sum of them

OpenStudy (anonymous):

Is that the answer?

OpenStudy (raden):

exactly, yeah... :)

OpenStudy (anonymous):

Okay, thank you. Do you mind helping me with another question? :o)

OpenStudy (raden):

cos4x + cos2x = 1 - 2sin^2 2x + 1 - 2sin^2 x = 2 - 2sin^2 2x - 2sin^2 x LHS = RHS

OpenStudy (raden):

i'll try

OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. \[f(x)=\frac{ x-7 }{ x+3 }\]\[g(x)=\frac{ -3x-7 }{ x-1 }\]

OpenStudy (anonymous):

@RadEn Do you know how to do this?

OpenStudy (raden):